OpenStudy (anonymous):

a - 1/2 = 3/5 please solve

OpenStudy (anonymous):

add 1/2 to each side, these are all the same that you've already posted. a = 11/10 or 1.1

OpenStudy (anonymous):

so your saying x+19=5 would equal 24

OpenStudy (anonymous):

Since you have a + then you need to subtract from both side. so that would be -14. Take the opposite function of what you have. So if there is a + then subtract from both sides, if there is a - add to both side.

OpenStudy (anonymous):

and what if theres -5.4=-1.5+h qould that work the same way and also could i get that answer

OpenStudy (anonymous):

Add 1.5 to both sides to solve for h. so you would get h = -3.9

OpenStudy (anonymous):

oh ok thanks for your help